# Modular degree and a conjecture of Watkins

### v2

@@ 3.5. Elliptic curves with trivial $2$-torsion
It is known from [18] that $E(\mathbb{Q})[2]$ is trivial for almost all the elliptic curves over $\mathbb{Q}$. Therefore, we are considering the most elliptic curves in this case. In this cases, we have $\mathrm{Gal}(\mathbb{Q}(E[2])/\mathbb{Q})$ is either $\mathbb{Z}/3\mathbb{Z}$, or symmetric group $S_{3}$. Note that $a(p)$ is even, if and only if, $\mathrm{Frob}_{p}$ has order $2$ in $S_{3}$. If the Galois group is $\mathbb{Z}/3\mathbb{Z}$, or equivalently, if $\Delta_{E}$ is a square, the set of such primes $p$ has a density exactly $\frac{1}{3}$. However, if the Galois group is $S_{3}$, the set of such primes $p$ has a density exactly $\frac{2}{3}$. With this observation, we prove the following.
@@
The proof of the second part follows immediately from (10). For the first part, denote $S_{E}$ as the set of all primes for which $a(p)$ is even. Viewing $\mathrm{Gal}(\mathbb{Q}(E[2])/\mathbb{Q})$ inside $S_{3}$, asking $a(p)$ to be even is equivalent to asking for $\mathrm{Frob}_{p}$ not to have order $3$. When the Galois group is $S_{3}$, or equivalently, if $\Delta_{E}$ is a square, the required density is $\frac{1}{3}$. Similarly in the $S_{3}$ extension case, the density is $\frac{2}{3}$. ∎
To deduce a consequence of Watkins’s conjecture, we study the rank of $E^{(D)}$ on an average. It is a conjecture of Goldfeld [17] that the average rank is $\frac{1}{2}$ for any elliptic curve $E/\mathbb{Q}$. Heath-brown [21] shows that BSD and GRH for the L-function of elliptic curves imply that the average is at most $\frac{3}{2}$. With this in hand, we have the following result.
Theorem 7.
@@
Denote $S$ as the set of squarefree integers $D$ for which Watkins’s conjecture is not true for $E^{(D)}$. It is enough to prove that $S$ has density 00. For the sake of contradiction, let us assume that the (lim-sup) density is $c>0$. It follows from Proposition 1 that, $\mathrm{rank}(E^{(D)})\gg_{E}\omega(D)$, for any square-free integer $D$. In particular, it follows from Heath-Brown’s result in [21] that
(13) $x\gg\sum_{\begin{subarray}{c}|D|\leq x,\\ D\in S\end{subarray}}\mathrm{rank}(E^{(D)})\gg_{E}\sum_{\begin{subarray}{c}|D|\leq x,\\ D\in S\end{subarray}}\omega(D).$
@@
$\sum_{\begin{array}[]{c}|D|\leq x,\\ D\in S\end{array}}\omega(|D|)=\sum_{k\leq\log x}kS_{k}(x)\geq cx\log x+O\left((\log x)^{2}\right),$
which is a contradiction to (13). ∎
In this regard, we also would like to point out that if $E[2](\mathbb{Q})$ is trivial, and $\nu_{2}(m_{E})=r$, then $\omega(N)\leq r$ provided that $E$ has non-split multiplication reduction at the prime dividing $N$. In that case, root number of $E/\mathbb{Q}$ is $-1$. If $r=1$, then $N$ is a prime power. In the prime case, note that Watkins’s conjecture follows from Kazalicki-Kohen in [24]. On the other hand, if $E$ has a split multiplicative reduction, $\omega(N)$ is either $1$ or $2$. Again, for the case $\omega(N)=1$, the conjecture is proved in [24] when $\Delta_{E}>0$. This leaves us with another possibility: what happens to the conjecture when $N=pq,E$ does not have split multiplicative at least at one prime and $E[2](\mathbb{Q})$ is trivial.
@@ 3.6. On a weaker question
Then how often $m_{E}$ is always even?
First of all, it follows from Lemma 1 that $\nu_{2}(m_{E})\geq 1$ for almost all elliptic curves over $\mathbb{Q}$ parameterized by $A,B\in\mathbb{Z}$. We now discuss the analogous distribution over any $1$-parameter family of elliptic curves. First, in the case of quadratic twists, it follows from Proposition 1 that $\nu_{2}(m_{E^{(D)}})\gg_{E}\omega(D)$. In particular, $m_{E^{(D)}}$ is even for almost all $D$. Now we ask the same for any $1$-parameter family of elliptic curves given by
@@
where $f(x),g(x)\in\mathbb{Q}[x]$ are two arbitrary polynomial. Then we have the following.
Theorem 8.
@@
#$\#\{t\in\mathbb{Z}:|t|\leq x,~{}\nu_{2}(m_{E_{t}})\geq 1\}=O\left(\frac{x}{\log x}\right).$
The proof of this theorem follows from an analogous version of Lemma 1, which we briefly state now.
Lemma 2.
@@
To prove this lemma, we follow a similar path as in the proof of Lemma 1. Consider the polynomial $\Delta(x)=\Delta_{E_{x}}\in\mathbb{Q}[x]$, and then it is enough to show that #$\#\{t\in\mathbb{Z}\mid|t|\leq x,\omega(\Delta(t))=1\}=O\left(\frac{x}{\log x}\right)$. For the Sieving argument, we consider $\{\mathcal{P}_{\lambda\vdash\mathrm{deg}(\Delta)}\}$, where $\mathrm{deg}(\Delta)$ denotes degree of the polynomial $\Delta$, and $\mathcal{P}_{\lambda}$ be the set of all primes $p$ for which $\Delta(x)\pmod{p}$ has a factorization of type $\lambda$ in $\mathbb{F}_{p}[x]$. Note that $\Delta(t)=0\pmod{p}$ for some prime $p$ if and only if, $p\in\mathcal{P}_{\lambda}$ for some partition $\lambda\vdash\mathrm{deg}(\Delta)$ such that $\lambda$ has at least one part which is $1$. Now for each such $\lambda$, we sieve modulo $\mathcal{P}_{\lambda}$.
Note that we got a significant advantage while working with the cases when $N_{E}$ has at least two prime factors. Now the question is, what happens for the families of elliptic curves whose conductors have only one prime factor? When $N$ is a prime, we know that $m_{E}$ is even when $\mathrm{rank}(E(\mathbb{Q}))>0$. For the rank 00 case, we have the following result.
Theorem 9.
@@
where the last equality follows from [25, Proposition 2.3]. In the case $p=7\pmod{12}$, we have $w_{i}=3$ for exactly one $i$. In this case, a similar argument applies. In the case $p=11\pmod{12}$, we have $w_{i}=2$ for exactly one $i$. In this case, $j(e_{i})=1728\in\mathbb{F}_{p}$. The proof then follows from Conjecture 2. ∎
Note that $m_{E}$ is not necessarily even when #$\#E_{\mathrm{tor}}$ is even, in other words, when $E(\mathbb{Q})[2]$ is non-trivial. In that case, according to Setzer’s classification, $p=17$ or of the form $u^{2}+64$. In the case, $p=17$, four non-isomorphic elliptic curves have LMFD levels respectively $17.a_{1},17.a_{2},17.a_{3}$ and $17.a_{4}$. It turns out that $m_{E}$ is even in the case of $17.a_{1},17.a_{2}$ and $17.a_{4}$. For primes of the form $p=u^{2}+64$, two non-isomorphic elliptic curves of conductor $p$ have level $p.a_{1}$ and $p.a_{2}$. In the case of $p.a_{1}$, the discriminant $p$ is positive and not a square. Then it follows from [24, Theorem 4] and arguing similarly as in the proof of [25, Theorem 1.4] we get $4|m_{E}\#E_{tor}$.
@@ 3.7. On the function field analogue
Let $E$ be an elliptic curve over $K$ of conductor $\mathfrak{n}_{E}=\mathfrak{n}\cdot\infty$ and it has multiplicative reduction at $\infty$. Then we have a result analogous to the modularity result, which asserts that parametrization map from the Drinfield modular curve $X_{0}(\mathfrak{n})$ to the elliptic curve $E$, $\phi_{E}:X_{0}(\mathfrak{n})\to E$ such that $\phi_{E}$ is non-trivial and of minimal possible degree [8, Theorem 2.1]. The degree of this map is defined to be the modular degree of $E$.
@@
Let $E$ be an elliptic curve with conductor $\mathfrak{n}_{E}=\mathfrak{n}\infty$, where $\mathfrak{n}$ is an ideal. Let $f_{E}$ be the primitive newform associated with $E$. We denote by ${\mathcal{W}}(\mathfrak{n})$ the 2-elementary abelian group of all Atkin-Lehner involutions. Let $f$ be a primitive newform; since $f$ is primitive, it is determined by its eigenvalues up to sign. We define ${\mathcal{W}}^{\prime}=\left\{W\in{\mathcal{W}}({\mathfrak{n}}):W(f_{E})=f_{E}\right\}$, which consists of the Hecke operators $W$ that preserve the newform $f_{E}$. Additionally, let $\kappa:=\dim_{\mathbb{F}_{2}}\left({\mathcal{W}}({\mathfrak{n}})/{\mathcal{W}}^{\prime}\right)+\dim_{\mathbb{F}_{2}}E(K)[2]$, where $E(K)[2]$ represents the $2$-torsion points of $E$ defined over $K$. Then, by [8, Proposition 3.2], we have the inequality $\omega_{K}({\mathfrak{n}})-\kappa\leq\nu_{2}(m_{E})$, where $\omega_{K}(n)$ denotes the number of prime divisors of ${\mathfrak{n}}$ in the field $K$, and $\nu_{2}(m_{E})$ represents the $2$-adic valuation of the modular degree $m_{E}$ associated with $E$. This is proved by realizing $M(\Gamma_{0}(\mathfrak{n}),\mathbb{Q}_{\ell})$ as the dual of $V_{\ell}(J_{0}(\mathfrak{(}n)))$. To be more precise, this identification shows that $\pi([W(D)])=\pi([D])$ for every divisor $D$ of degree 00 over $X^{0}(\mathfrak{n})$. The proof follows a similar path as that of [14, Proposition 2.1], as it assumes that the characteristic of $k$, denoted by $p$, is odd.
@@
In this functional field situation, we have the following rank bound given by Tate [39],
(14) $\mathrm{rank}_{\mathbb{Z}}(E(K))\leq\mathrm{deg}(\mathfrak{n}_{E})-4,$
@@
where $\mathfrak{n}_{E}$ is the finite part of the conductor of $E/K$. By utilizing the lower bound of $\nu_{2}(m_{E})$, Caro [8] potentially establishes the validity of the Watkins conjecture. In other words, denote $k^{\prime}$ as the splitting field of $\mathfrak{n}$ over the finite field $k$, and set $K^{\prime}=k^{\prime}(T)$. Then Watkins’s conjecture is true for the base change over $K^{\prime}$, i.e., for the elliptic curve $E^{\prime}=E\times_{\text{Spec }K}\text{Spec }K^{\prime}$, the Watkins’s conjecture is true. Here, $E$ represents a modular semi-stable elliptic curve defined over $K$ with conductor $\mathfrak{n}E=\mathfrak{n}{\infty}$. where $\mathfrak{n}$ is the conductor and $\mathfrak{n}_{\infty}$ is a finite field containing the splitting field of $\mathfrak{n}$ over $k$. Caro needed the base over $K^{\prime}$ to ensure that $\omega_{K^{\prime}}(\mathfrak{n})=\mathrm{deg}(\mathfrak{n})$, and the helps the lower bound of $\nu_{2}(m_{E})$ to beat the rank bound at (14). However, we can use the same technique to deduce the following, allowing us to work with a possibly smaller $K^{\prime}$.
Theorem 10.
@@
We omit the proof but note that $\kappa$ is always at most $3$; in that case, we have no improvement. However, $\kappa$ could still be $0,1$ or $2$, for instance, $\kappa=0$ if and only if, $E(K)[2]$ is trivial and $E$ has split-multiplicative reduction at every prime. In all these cases of $\kappa$, we have a possibly smaller extension of $K$ contained in $K^{\prime}$ that obeys the imposed condition at Theorem 10.
In the same article, Caro discusses Watkins’s conjecture for certain quadratic twists, i.e., twists by polynomials of even degree. The assumption on degree is needed to ensure that the twisted elliptic curve is modular, as remarked by Caro [8, Section 4]. With this, Caro showed that Watkins’s conjecture is true for any such even twist, provided that $\omega_{K}(g)\geq 3$ and $E/K$ are semi-stable. On the other hand, if $E/K$ is semi-stable but with non-trivial $E(K)[2]$, then similarly, as in the case of $\mathbb{Q}$, we also have the following function field analog of the rank bound
(15) $\mathrm{rank}_{\mathbb{Z}}(E(K))\leq\omega_{K}(\mathfrak{n}).$

### v3

@@ 3.5. Elliptic curves with trivial $2$-torsion
It is known from [grant00] that $E(\mathbb{Q})[2]$ is trivial for almost all the elliptic curves over $\mathbb{Q}$. Therefore, we are considering the most elliptic curves in this case. In this cases, we have $\mathrm{Gal}(\mathbb{Q}(E[2])/\mathbb{Q})$ is either $\mathbb{Z}/3\mathbb{Z}$, or symmetric group $S_{3}$. Note that $a(p)$ is even, if and only if, $\mathrm{Frob}_{p}$ has order $2$ in $S_{3}$. If the Galois group is $\mathbb{Z}/3\mathbb{Z}$, or equivalently, if $\Delta_{E}$ is a square, the set of such primes $p$ has a density exactly $\frac{1}{3}$. However, if the Galois group is $S_{3}$, the set of such primes $p$ has a density exactly $\frac{2}{3}$. With this observation, we prove the following.
@@
The proof of the second part follows immediately from (3.6). For the first part, denote $S_{E}$ as the set of all primes for which $a(p)$ is even. Viewing $\mathrm{Gal}(\mathbb{Q}(E[2])/\mathbb{Q})$ inside $S_{3}$, asking $a(p)$ to be even is equivalent to asking for $\mathrm{Frob}_{p}$ not to have order $3$. When the Galois group is $S_{3}$, or equivalently, if $\Delta_{E}$ is a square, the required density is $\frac{1}{3}$. Similarly in the $S_{3}$ extension case, the density is $\frac{2}{3}$. ∎
To deduce a consequence of Watkins’s conjecture, we study the rank of $E^{(D)}$ on an average. It is a conjecture of Goldfeld [goldfeld79] that the average rank is $\frac{1}{2}$ for any elliptic curve $E/\mathbb{Q}$. Heath-brown [HB2004] shows that BSD and GRH for the L-function of elliptic curves imply that the average is at most $\frac{3}{2}$. With this in hand, we have the following result.
Theorem 3.8.
@@
Denote $S$ as the set of squarefree integers $D$ for which Watkins’s conjecture is not true for $E^{(D)}$. It is enough to prove that $S$ has density 0. For the sake of contradiction, let us assume that the (lim-sup) density is $c>0$. It follows from Proposition 1 that, $\mathrm{rank}(E^{(D)})\gg_{E}\omega(D)$, for any square-free integer $D$. In particular, it follows from Heath-Brown’s result in [HB2004] that
(3.9) $x\gg\sum_{\begin{subarray}{c}|D|\leq x,\\ D\in S\end{subarray}}\mathrm{rank}(E^{(D)})\gg_{E}\sum_{\begin{subarray}{c}|D|\leq x,\\ D\in S\end{subarray}}\omega(D).$
@@
$\sum_{\begin{subarray}{c}|D|\leq x,\\ D\in S\end{subarray}}\omega(|D|)=\sum_{k\leq\log x}kS_{k}(x)\geq cx\log x+O\left((\log x)^{2}\right),$
which is a contradiction to (3.9). ∎
In this regard, we also would like to point out that if $E[2](\mathbb{Q})$ is trivial, and $\nu_{2}(m_{E})=r$, then $\omega(N)\leq r$ provided that $E$ has non-split multiplication reduction at the prime dividing $N$. In that case, root number of $E/\mathbb{Q}$ is 1-1. If $r=1$, then $N$ is a prime power. In the prime case, note that Watkins’s conjecture follows from Kazalicki-Kohen in [KK17singular]. On the other hand, if $E$ has a split multiplicative reduction, $\omega(N)$ is either $1$ or $2$. Again, for the case $\omega(N)=1$, the conjecture is proved in [KK17singular] when $\Delta_{E}>0$. This leaves us with another possibility: what happens to the conjecture when $N=pq,E$ does not have split multiplicative at least at one prime and $E[2](\mathbb{Q})$ is trivial.
@@ 3.6. On a weaker question
How often is $m_{E}$ always even?
First of all, it follows from Lemma 3.2 that $\nu_{2}(m_{E})\geq 1$ for almost all elliptic curves over $\mathbb{Q}$ parameterized by $A,B\in\mathbb{Z}$. We now discuss the analogous distribution over any $1$-parameter family of elliptic curves. First, in the case of quadratic twists, it follows from Proposition 1 that $\nu_{2}(m_{E^{(D)}})\gg_{E}\omega(D)$. In particular, $m_{E^{(D)}}$ is even for almost all $D$. Now we ask the same for any one-parameter family of elliptic curves given by
@@
where $f(x),g(x)\in\mathbb{Q}[x]$ are two arbitrary polynomials. Then we have the following.
Theorem 3.9.
@@
$\{t\in\mathbb{Z}:|t|\leq x,~\nu_{2}(m_{E_{t}})\geq 1\}=O\left(\frac{x}{\log x}\right).$
The proof of this theorem follows from an analogous version of Lemma 3.2, which we briefly state now.
Lemma 3.10.
@@
To prove this lemma, we follow a similar path as in the proof of Lemma 3.2. Consider the polynomial $\Delta(x)=\Delta_{E_{x}}\in\mathbb{Q}[x]$, and then it is enough to show that $\{t\in\mathbb{Z}\mid|t|\leq x,\omega(\Delta(t))=1\}=O\left(\frac{x}{\log x}\right)$. For the Sieving argument, we consider $\{\mathcal{P}_{\lambda\vdash\mathrm{deg}(\Delta)}\}$, where $\mathrm{deg}(\Delta)$ denotes degree of the polynomial $\Delta$, and $\mathcal{P}_{\lambda}$ be the set of all primes $p$ for which $\Delta(x)\pmod{p}$ has a factorization of type $\lambda$ in $\mathbb{F}_{p}[x]$. Note that $\Delta(t)=0\pmod{p}$ for some prime $p$ if and only if, $p\in\mathcal{P}_{\lambda}$ for some partition $\lambda\vdash\mathrm{deg}(\Delta)$ such that $\lambda$ has at least one part which is $1$. Now for each such $\lambda$, we sieve modulo $\mathcal{P}_{\lambda}$.
Note that we got a significant advantage while working with the cases when $N_{E}$ has at least two prime factors. Now the question is, what happens for the families of elliptic curves whose conductors have only one prime factor? When $N$ is a prime, we know that $m_{E}$ is even when $\mathrm{rank}(E(\mathbb{Q}))>0$. For the rank 0 case, we have the following result.
Theorem 3.11.
@@
where the last equality follows from [KK18watkins, Proposition 2.3]. In the case $p=7\pmod{12}$, we have $w_{i}=3$ for exactly one $i$. In this case, a similar argument applies. In the case $p=11\pmod{12}$, we have $w_{i}=2$ for exactly one $i$. In this case, $j(e_{i})=1728\in\mathbb{F}_{p}$. The proof then follows from Conjecture 2. ∎
Note that $m_{E}$ is not necessarily even when $E_{\mathrm{tor}}$ is even, in other words, when $E(\mathbb{Q})[2]$ is non-trivial. In that case, according to Setzer’s classification, $p=17$ or of the form $u^{2}+64$. In the case, $p=17$, four non-isomorphic elliptic curves have levels respectively $17.a_{1},17.a_{2},17.a_{3}$ and $17.a_{4}$. It turns out that $m_{E}$ is even in the case of $17.a_{1},17.a_{2}$ and $17.a_{4}$. For primes of the form $p=u^{2}+64$, two non-isomorphic elliptic curves of conductor $p$ have level $p.a_{1}$ and $p.a_{2}$. In the case of $p.a_{1}$, the discriminant $p$ is positive and not a square. Then it follows from [KK17singular, Theorem 4] and arguing similarly as in the proof of [KK18watkins, Theorem 1.4] we get $4|m_{E}\#E_{tor}$.
@@ 3.7. On the function field analogue
Let $E$ be an elliptic curve over $K$ of conductor $\mathfrak{n}_{E}=\mathfrak{n}\cdot\infty$ and it has multiplicative reduction at $\infty$. Then we have a result analogous to the modularity result, which asserts that parametrization map from the Drinfeld modular curve $X_{0}(\mathfrak{n})$ to the elliptic curve $E$, $\phi_{E}:X_{0}(\mathfrak{n})\to E$ such that $\phi_{E}$ is non-trivial and of minimal possible degree [caro2022watkins, Theorem 2.1]. The degree of this map is defined to be the modular degree of $E$.
@@
Let $E$ be an elliptic curve with conductor $\mathfrak{n}_{E}=\mathfrak{n}\infty$, where $\mathfrak{n}$ is an ideal. Let $f_{E}$ be the primitive newform associated with $E$. We denote by ${\mathcal{W}}(\mathfrak{n})$ the 2-elementary abelian group of all Atkin-Lehner involutions. Let $f$ be a primitive newform; since $f$ is primitive, it is determined by its eigenvalues up to sign. We define ${\mathcal{W}}^{\prime}=\left\{W\in{\mathcal{W}}({\mathfrak{n}}):W(f_{E})=f_{E}\right\}$, which consists of the Hecke operators $W$ that preserve the newform $f_{E}$. Additionally, let $\kappa:=\dim_{\mathbb{F}_{2}}\left({\mathcal{W}}({\mathfrak{n}})/{\mathcal{W}}^{\prime}\right)+\dim_{\mathbb{F}_{2}}E(K)[2]$, where $E(K)[2]$ represents the $2$-torsion points of $E$ defined over $K$. Then, by [caro2022watkins, Proposition 3.2], we have the inequality $\omega_{K}({\mathfrak{n}})-\kappa\leq\nu_{2}(m_{E})$, where $\omega_{K}(n)$ denotes the number of prime divisors of $\mathfrak{n}$} in the field $K$, and $\nu_{2}(m_{E})$ represents the $2$-adic valuation of the modular degree $m_{E}$ associated with $E$. This is proved by realizing $M(\Gamma_{0}(\mathfrak{n}),\mathbb{Q}_{\ell})$ as the dual of $V_{\ell}(J_{0}(\mathfrak{(}n)))$. To be more precise, this identification shows that $\pi([W(D)])=\pi([D])$ for every divisor $D$ of degree 0 over $X_{0}(\mathfrak{n})$. The proof follows a similar path as that of [DK13, Proposition 2.1], as it assumes that the characteristic of $k$, denoted by $p$, is odd.
@@
In this functional field situation, we have the following rank bound given by Tate [Tate95],
(3.10) $\mathrm{rank}_{\mathbb{Z}}(E(K))\leq\mathrm{deg}(\mathfrak{n}_{E})-4,$
@@
where $\mathfrak{n}_{E}$ is the finite part of the conductor of $E/K$. By utilizing the lower bound of $\nu_{2}(m_{E})$, Caro [caro2022watkins] potentially establishes the validity of the Watkins conjecture. In other words, denote $k^{\prime}$ as the splitting field of $\mathfrak{n}$ over the finite field $k$, and set $K^{\prime}=k^{\prime}(T)$. Then Watkins’s conjecture is true for the base change over $K^{\prime}$, i.e., for the elliptic curve $E^{\prime}=E\times_{\text{Spec }K}\text{Spec }K^{\prime}$, the Watkins’s conjecture is true. Here, $E$ represents a modular semi-stable elliptic curve defined over $K$ with conductor $\mathfrak{n}E=\mathfrak{n}{\infty}$. where $\mathfrak{n}$ is the conductor and $\mathfrak{n}_{\infty}$ is a finite field containing the splitting field of $\mathfrak{n}$ over $k$. Caro needed the base over $K^{\prime}$ to ensure that $\omega_{K^{\prime}}(\mathfrak{n})=\mathrm{deg}(\mathfrak{n})$, and the helps the lower bound of $\nu_{2}(m_{E})$ to beat the rank bound at (3.10). However, we can use the same technique to deduce the following, allowing us to work with a possibly smaller $K^{\prime}$.
Theorem 3.12.
@@
We omit the proof but note that $\kappa$ is always at most $3$; in that case, we have no improvement. However, $\kappa$ could still be $0,1$ or $2$, for instance, $\kappa=0$ if and only if, $E(K)[2]$ is trivial and $E$ has split-multiplicative reduction at every prime. In all these cases of $\kappa$, we have a possibly smaller extension of $K$ contained in $K^{\prime}$ that obeys the imposed condition at Theorem 3.12.
In the same article, Caro discusses Watkins’s conjecture for certain quadratic twists, i.e., twists by polynomials of even degree. The assumption on degree is needed to ensure that the twisted elliptic curve is modular, as remarked by Caro [caro2022watkins, Section 4]. With this, Caro showed that Watkins’s conjecture is true for any such even twist, provided that $\omega_{K}(g)\geq 3$ and $E/K$ are semi-stable. On the other hand, if $E/K$ is semi-stable but with non-trivial $E(K)[2]$, then similarly, as in the case of $\mathbb{Q}$, we also have the following function field analog of the rank bound
(3.11) $\mathrm{rank}_{\mathbb{Z}}(E(K))\leq\omega_{K}(\mathfrak{n}).$